OpenStudy (anonymous):

A country's population in 1993 was 127 million. In 2000 it was 132 million. Estimate the population in 2008 using the exponential growth formula. Round your answer to the nearest million. P = Aekt Now solve for k

OpenStudy (anonymous):

@Pizza7

OpenStudy (anonymous):

Use 1993 as year 0 and 127million as your starting point, thus your equation starts as P=127e^(k*0) since e^0 =1 that's 127milion just as it should be, next set it equal to 132mil and change the t to 7 years 132=127e^(k*7) now solve for k, first divide both sides by 127 132/127 = e^(k*7) next take the natural log of both sides leaving ln(132/127) = k*7 finally divide both sides by 7 k = (ln(132/127))/7 The result k= 0.005516405 rounded to 4 decimal places k= .0055 now put that in the equation and set the time to 15 years for 2008 P= 127e^(.0055*15) =137.9218 to the nearest million... 138 million people in 2008 hope that helps

OpenStudy (anonymous):

The result k= 0.005516405

OpenStudy (anonymous):

Yep thanks :)

OpenStudy (anonymous):

welcome